# Scalable Causal Structure Learning via Amortized Conditional Independence Testing

### v2

@@ 1 Introduction
Many scientific applications can be posed as a causal discovery problem where a directed acylic graph (DAG) is learned that models the causal dependencies within an observational dataset (e.g. Tennant et al. [2020], Ogburn et al. [2022]). In order to ensure the reliability of these findings, controlling the error rate on the set of causal discoveries is critical. Given finite data and noisy observations, exact determination of causal arrows in a DAG is impossible. As such, the causal discovery task is typically cast through the lens of statistical hypothesis testing for conditional independencies [Spirtes et al., 2000].
Although the problem of large-scale multiple hypothesis testing and uncertainty quantification is well studied in settings where purely associative relationships are the target discoveries [Benjamini and Hochberg, 1995, Goeman and Solari, 2014], large-scale causal learning produces a unique set of challenges. The number of possible DAGs scales superexponentially with the number of nodes in a graph, making causal learning an especially difficult problem when the number of variables in a dataset is large. Moreover, existing causal structure learning algorithms learn a graph with either no attempt to provide frequentist error guarantees [Chickering, 2003, Ramsey et al., 2017a, Zheng et al., 2018, Cundy et al., 2021b, Annadani et al., 2021, Cundy et al., 2021a] or require parametric assumptions on the data generating distribution which are unknown in practice [Strobl et al., 2019].
@@
In this paper, we introduce Scalable Causal Structure Learning (SCSL), a method for large-scale causal hypothesis testing that can scale to problems with hundreds of variables and thousands of potential edges for causal graphs with temporally separated sets of nodes. SCSL enables the use of black box machine learning models for hypothesis testing, requires no parametric assumptions, and returns a $p$-value for each edge under consideration. To scale to large graphs, SCSL recasts the causal search process as a discrete optimization problem. It then amortizes the dominant cost in causal structure identification: conditional independence testing over the combinatorial set of possible parent nodes in the causal graph. This avoids the combinatorial explosion in the candidate conditioning sets and reduces the search to a series of parallelizable optimization problems for each edge. We validate SCSL through semi-synthetic experiments using a cancer dataset that pairs genomic mutations at the primary tumor location of a cancer patient with information about metastases that have developed elsewhere in a patient’s body [Nguyen et al., 2022]. These simulation studies indicate that the method has high power, controls Type I error rate at the target level, and scales to larger graphs than existing methods.
@@
Classical algorithms for causal discovery like the SGS and PC algorithms (Spirtes et al. [2000]) convert causal structure learning into queries of conditional independence between nodes. The SGS algorithm draws an edge between two nodes if they are conditionally dependent given any possible subset of the remaining nodes. The PC algorithm simplifies this process by first deleting edges in the causal graph based on marginal independence testing. The size of the conditioning subsets is then allowed to increase by one for each subsequent round of CI testing, but fewer CI tests are required in each round as the algorithm only considers conditioning subsets of nodes that have remained adjacent to each other. The computational complexity of SGS matches the complexity of the PC in the worst case, but in most settings where the causal graph is reasonably sparse, the PC algorithm significantly reduces the number of independence tests needed to learn a causal graph. However, both methods are based on perfect knowledge of the CI structure and are computationally intractable beyond a few dozen nodes. Related work used the PC algorithm as a starting point and then relaxes assumptions around faithfulness (e.g. Ramsey et al. [2006]) or causal sufficiency (e.g. Spirtes [2001]), but these methods also rely on having perfect knowledge of the graph’s CI structure. The work of Strobl et al. [2019] extends the PC algorithm to generate edge-specific $p$-values, but only with provable Type I error control under the assumption of zero Type II error during the edge deletion stage of skeleton discovery.
@@
Other approaches to causal graph are score-based methods which maximize a score function such as BIC [D and Heckerman, 1997] or BDeau [Heckerman et al., 1995] over the space of all possible DAGs. Since the number of DAGs increases superexponentially with the number of nodes, approximate algorithms based on greedy search [Chickering, 2003, Ramsey et al., 2017a], coordinate descent [Aragam and Zhou, 2015, Fu and Zhou, 2013] or other methods are required. Other methods impose more stringent modeling assumptions such as linearity (e.g. Shimizu et al. [2011]) simplify the task. Follow-up work [Ramsey et al., 2017b, Solus et al., 2017] improves several of these methods to scale to datasets with thousands of nodes and potential edges using parallelization. However, none of these methods output edge specific $p$-values and are therefore not directly comparable to our proposed method.
@@
More recently, differentiable approaches to causal discovery have been proposed [Zheng et al., 2018], many of which enable Bayesian inference of the posterior distribution over DAGs [Cundy et al., 2021b, Annadani et al., 2021, Cundy et al., 2021a]. Unfortunately, these methods focus primarily on continuous rather than discrete data, limiting the application to datasets like the cancer dataset we consider in our simulations. Further, Bayesian uncertainty requires accurate model specification for valid posterior coverage. When the model is misspecified, Bayesian credible intervals can massively inflate Type I error. Similarly, Peters et al. [2016] weaken the faithfulness assumptions, but require repeated observations of variables across different intervention settings instead of relying on purely observational data. Overall, no method exists that is capable of providing Type I error control on individual edges across a large graph without making stringent parametric assumptions.
@@ 2 Methodology
We make the following assumptions about the graph and distribution of the data which are common in the literature [Spirtes et al., 2000].
@@
$\begin{split}&H_{0}:\exists S\subseteq\mathcal{Y}\setminus Y_{k}\text{ such that }X_{j}\perp\!\!\!\perp Y_{k}|S,X_{-j},\\ &H_{1}:X_{j}\text{ is not independent of }Y_{k}\text{ given }\{S,X_{-j}\}\text{ for all }S\subseteq Y_{-k}.\\ \end{split}$ (2)
@@ 2.1 Conditional Independence Testing
To test the null hypothesis that $X_{j}\perp\!\!\!\perp Y_{k}|S,X_{-j}$, we employ the generalized covariance measure (GCM) of Shah and Peters [2020]. The GCM recasts the problem of conditional independence into one about functional estimation of the expected conditional covariance
@@
and then tests whether this quantity is equal to 00. Any set of joint densities for $(\mathcal{X},\mathcal{Y})$ that are null will have this property, although it is possible for $X_{j}$ and $Y_{k}$ to be conditionally dependent but with 00 covariance. Therefore, this measure will only have power against the set of alternatives which have non-zero conditional covariance, which is likely to be the case in most practical settings.
The procedure constructs a test for the null hypothesis of 00 conditional covariance by first finding estimates $\hat{X}_{j}$ for the conditional expectation $\mathbb{E}[X_{j}|S,X_{-j}]$ and $\hat{Y}_{k}$ for the conditional expectation $\mathbb{E}[Y_{k}|S,X_{-j}]$. Any type of predictive modeling to construct this estimate is valid, for example construction of a neural net, so long as the product of the mean square errors of the two quantities is $o(n^{-1})$. Given $n$ samples $\{X_{j}^{i},Y_{k}^{i},S^{i},X_{-j}^{i}\}_{i=1}^{n}$, we denote $R_{i}=\left(X_{j}^{i}-\widehat{X}_{j}^{i}\right)\left(Y_{k}^{i}-\widehat{Y}_{k}^{i}\right)$, and the test statistic
@@
which will be distributed asymptotically as $\mathcal{N}(0,1)$ under the null, allowing us to construct a valid $p$-value. For a precise account of the technical conditions for the theorem, see Theorem 6 of Shah and Peters [2020] which we recall in the Appendix for completeness.
We note that our method is adaptable to other CI testing procedures such as conditional randomization tests (e.g. Candes et al. [2016], Tansey et al. [2022], Liu et al. [2021]). We focus on the GCM in this work because it requires less stringent modeling assumptions on the conditional distribution $X|Z$ and is less computationally burdensome.
@@ 2.2 Amortized Predictive Modeling
To make this process more computationally tractable, we will create a single predictive model for each element $Y_{k}\in\mathcal{Y}$ that has the flexibility to take in a choice of conditioning set $S\subseteq Y_{-k}$ as well as a choice of element $X_{j}\in\mathcal{X}$ and outputs an estimate of $\mathbb{E}[Y_{k}|X_{-j},S]$. We call this amortized predictive modeling because instead of creating bespoke models for each $S$, we localize the cost of training into a single flexible model. Formally, we wish to train a function for each $Y_{k}$ denoted as $\pi_{k}:\mathbb{R}^{|\mathcal{X}|}\times\mathbb{R}^{|\mathcal{Y}|}\times\{0,1\}^{|\mathcal{Y}|-1}\times[|\mathcal{X}|]\rightarrow\mathbb{R}$ where $[|\mathcal{X}|]:=\{1,...,|\mathcal{X}|\}$. The first two inputs into the function are the realized data points corresponding to $\mathcal{X}$ and $\mathcal{Y}$ while the second two inputs are user-specified masks which correspond to the set of nodes that the user wishes to include in the conditioning set. For the third input, which we label $Y^{\text{mask}}$, we interpret $S:=\{Y_{i}^{\text{mask}}\text{ s.t. }Y^{\text{mask}}_{i}=1\}$. For the fourth input, we can interpret the choice as corresponding to an element $X_{j}\in\mathcal{X}$ to not include in the conditioning set $X_{-j}$.
@@ 2.3 P-value optimization
In order to find the maximal $p$-value corresponding to Equation 3, one approach would be to exhaustively search over all possible subsets. However, this will not be computationally tractable for large graphs as the number of conditional independence tests needed to compute a $p$-value for a single edge scales exponentially with the number of nodes. The approach that we pursue is to learn the conditioning subset that minimizes the test statistic in Equation 5 through numerical optimization. However, since the conditioning set is a discrete rather than continuous variable, standard techniques such as gradient descent do not immediately apply. To overcome this, we use the the Gumbel-Softmax reparamaterization trick [Jang et al., 2017] to express the gradient with respect to the discrete variables we wish to optimize over with continuous relaxations.
@@ 3 Results
We perform a benchmark and simulation study centered on a $n=22,352$ dataset that pairs metastatic events with pre-metastasis tumor mutational info (Nguyen et al. [2022]). Previous studies have looked at the genomic landscape of metastases in different tissues such as brain [Brastianos et al., 2015]) and breast [Brown et al., 2017] metastases. However, this dataset is the only large database that pairs metastatic events with pre-metastasis tumor mutational info across a range of different sites in the body. In total, $234$ genes were sequenced for each patient along with $23$ secondary metastatic tissue sites (e.g. colon, breast, brain, etc.). Although records are collated across dozens of primary tumor site locations, we focus on the $10$ sites with the most patient records availability (breast, colon, liver, lung, ovary, pancreas, rectum, skin, and uterus) to ensure adequate sample size.
This dataset presents an opportunity for statistical models to discover new genomic biomarkers of metastatic potential. For each sequenced gene in a given tumor, we wish to know whether a mutation in that gene has a causal effect on metastasis to another site. Somatic mutations in genes are often highly correlated [Cheng et al., 2015], eliminating the use of simple association tests. Further, tumor colonization in one site may cause eventual metastasis to another site, such as liver metastases in colon cancer [Paschos et al., 2009]. It is therefore important to discover biomarkers with direct causal effect from the gene mutation to the specific metastatic site, such that intervention on that biomarker will have a positive effect on patient outcomes. Further, to ensure that discovered biomarkers are reliable, we wish to control the statistical error rate on reported causal links.
@@
Our primary baseline is the PC-p algorithm (Strobl et al. [2019]) as it is the only other existing methodology designed for frequentist error coverage. We use the same GCM test statistics described to perform CI tests for this method. We also employ the same simplifying assumptions used by SCSL to streamline the number of conditional independence tests that need to be evaluated — namely, by conditioning on all elements of $\mathcal{X}$ by default and orienting all edges between $\mathcal{X}$ and $\mathcal{Y}$ away from genetic sites and towards metastases.
@@ 3.3 Results on real dataset
Finally, we run the methodology on the real data. We stratify the dataset based on the primary tumor site location, and calculate $p$-values separately within each stratum to identify the secondary tumor location and gene combinations that are significant across different metastases. A rejection set is formed using the Benjamini-Hochberg (BH) [Benjamini and Hochberg, 1995] procedure with target FDR of 0.05 and is shown in Table 2. As a point of comparison, we also calculate the $p$-values for marginal tests of independence that the original paper used to identify connections. Figure 6 compares the marginal $p$-values used by the original paper with the causal $p$-values calculated through SCSL. Applying the BH procedure with the same target threshold results in $161$ rejections of marginal $p$-values. However, only $6$ of these $161$ remained in the causal rejection set.
The causal mutations discovered have mechanistic evidence in the biology literature. For instance, CDH1 has been mechanistically investigated in breast cancer models of metastasis through its connection to asparagine [Knott et al., 2018]. Colon cancers are often treated with EGFR-inhibitors, which are ineffective in the presence of KRAS mutants which continue to activate the MAPK pathway, leading to eventual metastasis [Prenen et al., 2010]. TP53 mutations in certain pancreas cancers have been shown to increase fibrosis, enabling tumors to better evade the immune system and increasing metastatic potential [Maddalena et al., 2021]. While these mechanistic links between the primary site, gene, and general metastasis are known, the site-specific patterns have not been investigated. Thus, our causal testing results may provide valuable guidance to scientists and clinicians considering the utility of invasive patient monitoring.
Figure 6: Plot of the $161$ discoveries rejected using the BH procedure with target FDR of $0.05$ on the marginal independence associations. Only $6$ discoveries remain when applying the BH procedure with the same target threshold on the causal $p$-values.
@@
Table 2: List of all $6$ genes and tumor combinations (corresponding to red markers on right graph) identified as significant at each primary site after forming a rejection set using the Benhamini-Hochberg procedure with target FDR of $0.05$. The causal $p$-values are significantly more conservative, leading to fewer rejections.

### v3

@@ 1 Introduction
Many scientific applications can be posed as a causal discovery problem where a directed acylic graph (DAG) is learned that models the causal dependencies within an observational dataset (e.g. application1, application2). In order to ensure the reliability of these findings, controlling the error rate on the set of causal discoveries is critical. Given finite data and noisy observations, exact determination of causal arrows in a DAG is impossible. As such, the causal discovery task is typically cast through the lens of statistical hypothesis testing for conditional independencies [spirtes:etal:2000:pc-algorithm].
Although the problem of large-scale multiple hypothesis testing and uncertainty quantification is well studied in settings where purely associative relationships are the target discoveries [bh_procedure, goeman_hyptesting], large-scale causal learning produces a unique set of challenges. The number of possible DAGs scales superexponentially with the number of nodes in a graph, making causal learning an especially difficult problem when the number of variables in a dataset is large. Moreover, existing causal structure learning algorithms learn a graph with either no attempt to provide frequentist error guarantees [GES_orig, Ramsey2017AMV, zheng2018dags, cundy2021bcd, Annadani2021VariationalCN, BCD_NETS] or require parametric assumptions on the data generating distribution which are unknown in practice [strobl:etal:2016:pc-p-algorithm].
@@
In this paper, we introduce Scalable Causal Structure Learning (SCSL), a method for large-scale causal hypothesis testing that can scale to problems with hundreds of variables and thousands of potential edges for causal graphs with temporally separated sets of nodes. SCSL enables the use of black box machine learning models for hypothesis testing, requires no parametric assumptions, and returns a $p$-value for each edge under consideration. To scale to large graphs, SCSL recasts the causal search process as a discrete optimization problem. It then amortizes the dominant cost in causal structure identification: conditional independence testing over the combinatorial set of possible parent nodes in the causal graph. This avoids the combinatorial explosion in the candidate conditioning sets and reduces the search to a series of parallelizable optimization problems for each edge. We validate SCSL through semi-synthetic experiments using a cancer dataset that pairs genomic mutations at the primary tumor location of a cancer patient with information about metastases that have developed elsewhere in a patient’s body [data_source]. These simulation studies indicate that the method has high power, controls Type I error rate at the target level, and scales to larger graphs than existing methods.
@@
Classical algorithms for causal discovery like the SGS and PC algorithms (spirtes:etal:2000:pc-algorithm) convert causal structure learning into queries of conditional independence between nodes. The SGS algorithm draws an edge between two nodes if they are conditionally dependent given any possible subset of the remaining nodes. The PC algorithm simplifies this process by first deleting edges in the causal graph based on marginal independence testing. The size of the conditioning subsets is then allowed to increase by one for each subsequent round of CI testing, but fewer CI tests are required in each round as the algorithm only considers conditioning subsets of nodes that have remained adjacent to each other. The computational complexity of SGS matches the complexity of the PC in the worst case, but in most settings where the causal graph is reasonably sparse, the PC algorithm significantly reduces the number of independence tests needed to learn a causal graph. However, both methods are based on perfect knowledge of the CI structure and are computationally intractable beyond a few dozen nodes. Related work used the PC algorithm as a starting point and then relaxes assumptions around faithfulness (e.g. ramsey_conservative_pc) or causal sufficiency (e.g. FCI), but these methods also rely on having perfect knowledge of the graph’s CI structure. The work of strobl:etal:2016:pc-p-algorithm extends the PC algorithm to generate edge-specific $p$-values, but only with provable Type I error control under the assumption of zero Type II error during the edge deletion stage of skeleton discovery.
@@
Other approaches to causal graph are score-based methods which maximize a score function such as BIC [BIC] or BDeau [bdeau] over the space of all possible DAGs. Since the number of DAGs increases superexponentially with the number of nodes, approximate algorithms based on greedy search [GES_orig, Ramsey2017AMV], coordinate descent [coord_aragam, coord_qing] or other methods are required. Other methods impose more stringent modeling assumptions such as linearity (e.g. LINGAMdirect) simplify the task. Follow-up work [fges, BOSS] improves several of these methods to scale to datasets with thousands of nodes and potential edges using parallelization. However, none of these methods output edge specific $p$-values and are therefore not directly comparable to our proposed method.
@@
More recently, differentiable approaches to causal discovery have been proposed [zheng2018dags], many of which enable Bayesian inference of the posterior distribution over DAGs [cundy2021bcd, Annadani2021VariationalCN, BCD_NETS]. Unfortunately, these methods focus primarily on continuous rather than discrete data, limiting the application to datasets like the cancer dataset we consider in our simulations. Further, Bayesian uncertainty requires accurate model specification for valid posterior coverage. When the model is misspecified, Bayesian credible intervals can massively inflate Type I error. Similarly, icp_algo weaken the faithfulness assumptions, but require repeated observations of variables across different intervention settings instead of relying on purely observational data. Overall, no method exists that is capable of providing Type I error control on individual edges across a large graph without making stringent parametric assumptions.
@@ 2 Methodology
We make the following assumptions about the graph and distribution of the data which are common in the literature [spirtes:etal:2000:pc-algorithm].
@@
[…]
@@ 2.1 Conditional Independence Testing
To test the null hypothesis that $X_{j}\perp\!\!\!\perp Y_{k}|S,X_{-j}$, we employ the generalized covariance measure (GCM) of hardness_CI. The GCM recasts the problem of conditional independence into one about functional estimation of the expected conditional covariance
@@
and then tests whether this quantity is equal to 0. Any set of joint densities for $(\mathcal{X},\mathcal{Y})$ that are null will have this property, although it is possible for $X_{j}$ and $Y_{k}$ to be conditionally dependent but with 0 covariance. Therefore, this measure will only have power against the set of alternatives which have non-zero conditional covariance, which is likely to be the case in most practical settings.
The procedure constructs a test for the null hypothesis of 0 conditional covariance by first finding estimates $\hat{X}_{j}$ for the conditional expectation $\mathbb{E}[X_{j}|S,X_{-j}]$ and $\hat{Y}_{k}$ for the conditional expectation $\mathbb{E}[Y_{k}|S,X_{-j}]$. Any type of predictive modeling to construct this estimate is valid, for example construction of a neural net, so long as the product of the mean square errors of the two quantities is $o(n^{-1})$. Given $n$ samples $\{X_{j}^{i},Y_{k}^{i},S^{i},X_{-j}^{i}\}_{i=1}^{n}$, we denote $R_{i}=\left(X_{j}^{i}-\widehat{X}_{j}^{i}\right)\left(Y_{k}^{i}-\widehat{Y}_{k}^{i}\right)$, and the test statistic
@@
which will be distributed asymptotically as $\mathcal{N}(0,1)$ under the null, allowing us to construct a valid $p$-value. For a precise account of the technical conditions for the theorem, see Theorem 6 of hardness_CI which we recall in the Appendix for completeness.
We note that our method is adaptable to other CI testing procedures such as conditional randomization tests (e.g. model_X, tansey_HRT, distilled_CRT). We focus on the GCM in this work because it requires less stringent modeling assumptions on the conditional distribution $X|Z$ and is less computationally burdensome.
@@ 2.2 Amortized Predictive Modeling
To make this process more computationally tractable, we will create a single predictive model for each element $Y_{k}\in\mathcal{Y}$ that has the flexibility to take in a choice of conditioning set $S\subseteq Y_{-k}$ as well as a choice of element $X_{j}\in\mathcal{X}$ and outputs an estimate of $\mathbb{E}[Y_{k}|X_{-j},S]$. We call this amortized predictive modeling because instead of creating bespoke models for each $S$, we localize the cost of training into a single flexible model. Formally, we wish to train a function for each $Y_{k}$ denoted as $\pi_{k}:\mathbb{R}^{|\mathcal{X}|}\times\mathbb{R}^{|\mathcal{Y}|}\times\{0,1\}^{|\mathcal{Y}|-1}\times[|\mathcal{X}|]\rightarrow\mathbb{R}$ where $|\mathcal{X}|]:=\{1,...,|\mathcal{X}|\}$. The first two inputs into the function are the realized data points corresponding to $\mathcal{X}$ and $\mathcal{Y}$ while the second two inputs are user-specified masks which correspond to the set of nodes that the user wishes to include in the conditioning set. For the third input, which we label $Y^{\text{mask}}$, we interpret $S:=\{Y_{i}^{\text{mask}}\text{ s.t. }Y^{\text{mask}}_{i}=1\}$. For the fourth input, we can interpret the choice as corresponding to an element $X_{j}\in\mathcal{X}$ to not include in the conditioning set $X_{-j}$.
@@ 2.3 P-value optimization
In order to find the maximal $p$-value corresponding to Equation 3, one approach would be to exhaustively search over all possible subsets. However, this will not be computationally tractable for large graphs as the number of conditional independence tests needed to compute a $p$-value for a single edge scales exponentially with the number of nodes. Building on recent work in differentiable casual structure learning [brouillard2020differentiable, kalainathan2022structural, ng2022masked], the approach we pursue is to learn the conditioning subset that minimizes the test statistic in Equation 5 through numerical optimization. Since the conditioning set is a discrete rather than continuous variable, standard techniques such as gradient descent do not immediately apply. To overcome this, we use the Gumbel-Softmax reparamaterization trick [gumbel_softmax] to express the gradient with respect to the discrete variables we wish to optimize with continuous relaxations.
@@ 3 Results
We perform a benchmark and simulation study centered on a $n=22,352$ dataset that pairs metastatic events with pre-metastasis tumor mutational info (data_source). Previous studies have looked at the genomic landscape of metastases in different tissues such as brain [brastianos2015genomic]) and breast [brown2017phylogenetic] metastases. However, this dataset is the only large database that pairs metastatic events with pre-metastasis tumor mutational info across a range of different sites in the body. In total, $234$ genes were sequenced for each patient along with $23$ secondary metastatic tissue sites (e.g. colon, breast, brain, etc.). Although records are collated across dozens of primary tumor site locations, we focus on the $10$ sites with the most patient records availability (breast, colon, liver, lung, ovary, pancreas, rectum, skin, and uterus) to ensure adequate sample size.
This dataset presents an opportunity for statistical models to discover new genomic biomarkers of metastatic potential. For each sequenced gene in a given tumor, we wish to know whether a mutation in that gene has a causal effect on metastasis to another site. Somatic mutations in genes are often highly correlated [cheng2015memorial], eliminating the use of simple association tests. Further, tumor colonization in one site may cause eventual metastasis to another site, such as liver metastases in colon cancer [paschos2009role]. It is therefore important to discover biomarkers with direct causal effect from the gene mutation to the specific metastatic site, such that intervention on that biomarker will have a positive effect on patient outcomes. Further, to ensure that discovered biomarkers are reliable, we wish to control the statistical error rate on reported causal links.
@@
Our primary baseline is the PC-p algorithm (strobl:etal:2016:pc-p-algorithm) as it is the only other existing methodology designed for frequentist error coverage. We use the same GCM test statistics described to perform CI tests for this method. We also employ the same simplifying assumptions used by SCSL to streamline the number of conditional independence tests that need to be evaluated — namely, by conditioning on all elements of $\mathcal{X}$ by default and orienting all edges between $\mathcal{X}$ and $\mathcal{Y}$ away from genetic sites and towards metastases.
@@ 3.3 Results on real dataset
Finally, we run the methodology on the real data. We stratify the dataset based on the primary tumor site location, and calculate $p$-values separately within each stratum to identify the secondary tumor location and gene combinations that are significant across different metastases. A rejection set is formed using the Benjamini-Hochberg (BH) [bh_procedure] procedure with target FDR of 0.05 and is shown in Table 2. As a point of comparison, we also calculate the $p$-values for marginal tests of independence that the original paper used to identify connections. Figure 6 compares the marginal $p$-values used by the original paper with the causal $p$-values calculated through SCSL. Applying the BH procedure with the same target threshold results in $161$ rejections of marginal $p$-values. However, only $6$ of these $161$ remained in the causal rejection set.
The causal mutations discovered have mechanistic evidence in the biology literature. For instance, CDH1 has been mechanistically investigated in breast cancer models of metastasis through its connection to asparagine [knott2018asparagine]. Colon cancers are often treated with EGFR-inhibitors, which are ineffective in the presence of KRAS mutants which continue to activate the MAPK pathway, leading to eventual metastasis [prenen2010new]. TP53 mutations in certain pancreas cancers have been shown to increase fibrosis, enabling tumors to better evade the immune system and increasing metastatic potential [maddalena:etal:2021:pdac-tp53-mets]. While these mechanistic links between the primary site, gene, and general metastasis are known, the site-specific patterns have not been investigated. Thus, our causal testing results may provide valuable guidance to scientists and clinicians considering the utility of invasive patient monitoring.
Figure 6: Plot of the $161$ discoveries rejected using the BH procedure with target FDR of $0.05$ on the marginal independence associations. Only 66 discoveries remain when applying the BH procedure with the same target threshold on the causal $p$-values.
@@
Table 2: List of all 66 genes and tumor combinations (corresponding to red markers on right graph) identified as significant at each primary site after forming a rejection set using the Benhamini-Hochberg procedure with target FDR of $0.05$. The causal $p$-values are significantly more conservative, leading to fewer rejections.